OpenStudy (anonymous):

I don't get this: Walter had 3/5 yard of wrapping paper. He used 3/4 of it. How much wrapping paper did he use? I found the common denominator and I get a subtraction problem of 12/20 - 15/20, which is obviously wrong. The answer is, according to my summer school teacher, 9/20. ???

OpenStudy (anonymous):

yes its 9/20ths because you have to take 3/4 of 3/5ths. 3/5ths is .6. take .6 multiply by 3/4 and you get .45 which is 9/20ths.

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 }\times \frac{ 3 }{ 4 }=\frac{ 9 }{ 20 }\]